find the correct value of x so that(x,-4) is a solution to 5x-2y=33
Answer by ikleyn(52932)

.
Simply substitute y= -4 into the equation and then solve it for x:


    5x - 2*(-4) = 33

    5x + 8 = 33

    5x = 33 - 8 = 25  ====>  x = 25/5 = 5.    ANSWER

Solved.